# LinGen: Towards High-Resolution Minute-Length Text-to-Video Generation with Linear Computational Complexity

### v1

@@ 1 Introduction
To address the above challenge, we propose a Linear-complexity text-to-video Generation (LinGen) framework that scales linearly in the number of pixels in generated videos. To the best of our knowledge, LinGen is the first to enable photorealistic high-resolution minute-length video generation at a high frame rate on a single GPU without video extension, super-resolution, or compromising quality. It not only addresses the aforementioned adjacency preservation issue, but also comprehensively enhances the short-, medium-, and long-range correlations while maintaining linear complexity. LinGen replaces the self-attention layers in DiTs with our proposed linear-complexity MATE blocks. Each MATE block is composed of an MA-branch and a TE-branch. The MA-branch consists of a bidirectional Mamba2 [6] (a transformer-format SSM variant) block equipped with our proposed Rotary-Major Scan (RMS) and review tokens. RMS rearranges 3D token tensors in the latent space before they enter the bidirectional Mamba2 block, enhancing short-range correlations. To alleviate the inherent long-range correlation decay of SSMs, review tokens provide an overview of the processed token sequences to the hidden state of Mamba2 blocks at the start of sequence processing, to calibrate long-range correlations. The TE-branch is a novel TEmporal Swin Attention (TESA) block. It computes correlations among short-range spatially adjacent and medium-range temporally adjacent tokens, focusing on addressing the adjacency preservation issue and improving video consistency. Note that LinGen is orthogonal to sampling distillation and can potentially be combined with it to further boost its efficiency. Our contributions can be summarized as follows.
@@ 4.3 Comparing Quality to State-of-the-Art Models
Class Objects Action Relatio. Style Style Consist. Score
@@ 4.5 Ablation Experiments
Model Latency/s
LinGen (default setting) 102
LinGen w/o TESA 94 (-8)
LinGen w/o RMS 99 (-3)
LinGen w/ Zigzag 144 (+42)
LinGen w/ Mamba 127 (+25)
LinGen w/o MA-branch 65 (-37)
LinGen w/o review tokens 98 (-4)
@@ Acknowledgment
Appendix A Visual Examples
We have posted visual examples of videos generated using our LinGen model and other baseline models at a local website. You can open the index.html file with your website browser to investigate them. This local website is based on an online website template [39]. These visual examples include:
@@ Appendix A Visual Examples
  • Comparisons with existing video generation works. Our baselines are typical open-source models (see Fig. 14), including T2V-Turbo [28], CogVideoX-5B [71], and OpenSora v1.2 [75], state-of-the-art accessible commercial models (see Fig. 15), including Kling [24], Runway Gen3 [47], and LumaLabs [33], and minute-length video generation trials (see Fig. 16), including Loong [65] and PA-VDM [66]. Note that PA-VDM has not yet released the code and prompts. Thus, we selected one LinGen-generated video similar to their demo video for reference.
@@
Appendix B Comparisons with Prior Works
In this section, we first supplement VBench results reported in Sec. B.1 in order to compare with more models and discuss the limitations of VBench. Then, we present visual examples of the generated videos to provide comparisons with prior works and include additional human evaluation results in Sec. B.2 to demonstrate high quality of videos generated by LinGen.
B.1 Automatic Metrics: VBench Results
@@
B.2 Visual Examples and Human Evaluation
@@
Appendix C More Ablation Experiments
@@
Appendix D Model Implementation Details
In this section, we first provide more details of our model backbone in Sec. D.1. Then, we compare Mamba and Mamba2 and present their technical details in Sec. D.2. Finally, we give the details of our training recipe in Sec. D.3.
D.1 Backbone Details
@@
D.2 Mamba and Mamba2
SSMs have gained popularity in the field of natural language processing due to their high efficiency and strong performance in handling long sequences [14, 13]. Mamba [12], as a variant of SSM, enhances efficiency significantly by incorporating dynamic parameters into the SSM structure and developing algorithms optimized for better hardware compatibility. Based on this, Mamba2 [6] unifies SSMs and masks efficient attention by proposing a special SSM with an attention format (i.e., Structured State Space Duality). Mamba2 removes sequential linear projections that are used in Mamba and produces SSM parameters $A,B,C$ in parallel. The normalization layer in Mamba2 is the same as that in [51]. It improves stability. As mentioned in our main paper, the FLOPs cost of a bidirectional Mamba2 module is given by
@@ D.2 Mamba and Mamba2
D.3 Training Recipe Details
In this Section, we introduce our progressive training recipe in Sec. D.3.1. Then, we discuss our text-to-image and text-to-video hybrid training setting in Sec. D.3.2. We describe the details of our training datasets and quality-tuning design in Sec. D.3.3.
@@ D.3 Training Recipe Details
D.3.1 Progressive Training Recipe
@@ D.3.2 Hybrid Training
D.3.3 Quality Tuning and Datasets

### v2

@@ 1 Introduction
To address the above challenge, we propose a Linear-complexity text-to-video Generation (LinGen) framework that scales linearly in the number of pixels in generated videos. To the best of our knowledge, LinGen is the first to enable photorealistic high-resolution minute-length video generation at a high frame rate on a single GPU without video extension, super-resolution, or compromising quality. It not only addresses the aforementioned adjacency preservation issue (see Supplementary Sec. A), but also comprehensively enhances the short-, medium-, and long-range correlations while maintaining linear complexity. LinGen replaces the self-attention layers in DiTs with our proposed linear-complexity MATE blocks. Each MATE block is composed of an MA-branch and a TE-branch. The MA-branch consists of a bidirectional Mamba2 [6] (a transformer-format SSM variant) block equipped with our proposed Rotary-Major Scan (RMS) and review tokens. RMS rearranges 3D token tensors in the latent space before they enter the bidirectional Mamba2 block, enhancing short-range correlations. To alleviate the inherent long-range correlation decay of SSMs, review tokens provide an overview of the processed token sequences to the hidden state of Mamba2 blocks at the start of sequence processing, to calibrate long-range correlations. The TE-branch is a novel TEmporal Swin Attention (TESA) block. It computes correlations among short-range spatially adjacent and medium-range temporally adjacent tokens, focusing on addressing the adjacency preservation issue and improving video consistency. Note that LinGen is orthogonal to sampling distillation and can potentially be combined with it to further boost its efficiency. Our contributions can be summarized as follows.
@@ 4.3 Comparing Quality to State-of-the-Art Models
Class Objects Action Relatio. Style Style Consist. Score
@@ 4.5 Ablation Experiments
TE-branch MA-branch Latency/s
Scan Method Mamba version Review Tokens
RMS Mamba2 102
RMS Mamba2 94 (-8)
Mamba2 99 (-3)
Zigzag Mamba2 144 (+42)
RMS Mamba 127 (+25)
RMS Mamba2 98 (-4)
65 (-37)
@@ Acknowledgment
Appendix A Adjacency Preservation
Vanilla Mamba2 cannot be scaled to process huge images and video tokens well due to its long-range decay and the well-known adjacency preservation issue (see Sec. 1 of the main paper), causing distorted and inconsistent videos. Loss comparisons in Fig. 11 of the main paper and ablative videos (see Sec. B) validate the effectiveness of RMS and TESA. Mamba models scan image and video tokens into a sequence, where the minimum distance between originally adjacent tokens in $k$ layers reflects their most precise correlation. For an $H\times W\times T$ token tensor, we compute its average, $d_{k}$, among adjacent tokens in a $2\times 2\times 2$ cube and plot $d_{k}$ in Fig. 13 for $H=W=T=32$. RMS achieves the same $d_{k}$ as Zigzag while being much more efficient and scalable (see Table 2 of the main paper). RMS and TESA thoroughly address the adjacency preservation issue.
Figure 13: Average minimum distance between adjacent tokens.
Appendix B Visual Examples
We provide visual examples next. They can also be found on our project website.
@@ Appendix B Visual Examples
  • Comparisons with existing video generation works. Our baselines are typical open-source models (see Fig. 15), including T2V-Turbo-v2 [28], CogVideoX-5B [72], and OpenSora v1.2 [76], state-of-the-art accessible commercial models (see Fig. 16), including Kling [24], Runway Gen3 [46], and LumaLabs [33], and minute-length video generation trials (see Fig. 17), including Loong [65] and PA-VDM [66]. Note that PA-VDM has not yet released the code and prompts. Thus, we selected one LinGen-generated video similar to their demo video for reference.
@@
Appendix C Comparisons with Prior Works
In this section, we first supplement VBench results reported in Sec. C.1 in order to compare with more models and discuss the limitations of VBench. Then, we present visual examples of the generated videos to provide comparisons with prior works and include additional human evaluation results in Sec. C.2 to demonstrate high quality of videos generated by LinGen.
C.1 Automatic Metrics: VBench Results
@@
C.2 Visual Examples and Human Evaluation
@@
Appendix D More Ablation Experiments
@@
Appendix E Model Implementation Details
In this section, we first provide more details of our model backbone in Sec. E.1. Then, we compare Mamba and Mamba2 and present their technical details in Sec. E.2. Finally, we give the details of our training recipe in Sec. E.3.
E.1 Backbone Details
@@
E.2 Mamba and Mamba2
SSMs have gained popularity in the field of natural language processing due to their high efficiency and strong performance in handling long sequences [14, 13]. Mamba [12], as a variant of SSM, enhances efficiency significantly by incorporating dynamic parameters into the SSM structure and developing algorithms optimized for better hardware compatibility. Early explorations [57, 69] replaced the attention layers in diffusion models with SSMs, such as Mamba, to perform image generation, but these prototypes stayed relatively small. To unlock better efficiency at large scale, Mamba2 [6] unifies SSMs and the masked efficient attention by proposing a special SSM with an attention format (i.e., Structured State Space Duality). Mamba2 removes sequential linear projections that are used in Mamba and produces SSM parameters $A,B,C$ in parallel. The normalization layer in Mamba2 is the same as that in [50]. It improves stability. As mentioned in our main paper, the FLOPs cost of a bidirectional Mamba2 module is given by
@@ E.2 Mamba and Mamba2
E.3 Training Recipe Details
In this Section, we introduce our progressive training recipe in Sec. E.3.1. Then, we discuss our text-to-image and text-to-video hybrid training setting in Sec. E.3.2. We describe the details of our training datasets and quality-tuning design in Sec. E.3.3.
@@ E.3 Training Recipe Details
E.3.1 Progressive Training Recipe
@@ E.3.2 Hybrid Training
E.3.3 Quality Tuning and Datasets